OpenStudy (anonymous):

Integral of (1/t)(cos3t)

OpenStudy (dumbcow):

sorry i left you there... ok so cos(3t) is in numerator and t in denominator ? in that case: you have to use integration by parts a couple times \[\int\limits_{}^{}u *dv = uv - \int\limits_{}^{} v *du\] let u = cos(3t) dv = 1/t du = -3sin(3t) v = ln(t)

OpenStudy (dumbcow):

i apologize again, i didn't work all the way through the problem .... doing it this way leads nowhere this indefinite integral can't be done algebraically http://www.wolframalpha.com/input/?i=integrate+cos%283t%29%2Ft+dt